OpenStudy (xmando98):

Consider the following binomial experiment: A survey shows 56% of households in Centercity own a DVD player. In a random sample of 19 households in this city, what is the probability that exactly 13 households own a DVD player? a) 0.1052 b) 0.1051 c) 0.1053 d) 0.1049 e) 0.1050 f) None of the above.

OpenStudy (xmando98):

can somebody help me out with this please? I did it but ended up getting a completely different number

OpenStudy (xmando98):

I got this 0.00033481966

OpenStudy (xmando98):

i guessed the correct answer which was d) .1049

OpenStudy (mathmale):

I'd suggest you share your work. Seeing how you got your result helps the helper more than just seeing your proposed answer. In this binomial probability problem, the probability that a family owns a DVD player is p=0.56. The number of samples is n=19. You could 1) calculate the relevant probability by hand from the appropriate formula, or 2) use a calculator that has statistics capabilities built in, or 3) use a table of binomial probabilities. Which are you going to try?

OpenStudy (xmando98):

Thank you for time anyway :)

OpenStudy (xmando98):

*Thank you for your time anyway :)

OpenStudy (eliesaab):

Apply this formula for n=19, k=13 and p=.56 \[ \frac{n!}{k!(n-k!}p^k (1-p)^{n-k} \]

OpenStudy (eliesaab):

If you do that, you get \[\frac {19!}{13! 6!}0.56^{13} (1-0.56)^6 =0.104868 \] What should be your answer